OpenStudy (anonymous):

find two numbers whose sum is 30 such that the sum of their squares is minimum

OpenStudy (cwrw238):

x + y = 30 S = x^2 + y^2 - find minimum S

OpenStudy (cwrw238):

we can do this using differentiation from the first equation y = 30 -x so S = x^2 + (30 - x)^2

OpenStudy (cwrw238):

can you find the derivative dS / dx?

OpenStudy (cwrw238):

or haven't you done any calculus?

OpenStudy (anonymous):

is that 4x-60?

OpenStudy (anonymous):

from that we can get x=15 ..is that right?

OpenStudy (cwrw238):

dS/dx = 4x - 60 x = 15 yes

OpenStudy (anonymous):

thank you

OpenStudy (cwrw238):

so both x and y are 15

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

to check if it is a minimum find the second derivative which is 4 positive - therefore minimum

OpenStudy (anonymous):

ok thanks alot sir